Adam wants tomake a 52% acid solution. He has already poured 2 fluid ounces of a 55% acid solution into a beaker. How many fluid ounces of a 50% acid solution must he add to make this mix.
Answer by lwsshak3(11628)

Adam wants tomake a 52% acid solution. He has already poured 2 fluid ounces of a 55% acid solution into a beaker. How many fluid ounces of a 50% acid solution must he add to make this mix.
**
let x=fluid ounces of 50% solution must be added to make desired mix
55%*2+50%x=52%(2+x)
.55*2+.5x=1.04+.52x
.02x=1.10-10.4=.06
.02x=.06
x=3 fluid ounces
ans:
3 fluid ounces of 50% solution must be added to make desired mix